# Population Protocols Revisited: Parity and Beyond

### v2

@@ 3.6 Proof of Theorem 1
By Theorem 2, the fast MC+ parity protocol can be combined with a slow parity protocol, resulting in an efficient, stable parity protocol. This protocol is silent because the slow parity protocol is silent. The total number of rounds of the MC+ parity protocol is $R=O(\log^{2}n)$. By Lemma 2, the slow parity protocol requires polynomial time. By Theorem 2, the expected running time of the fast stable parity protocol is $O(\log^{3}n)$. The total number of states in the combined parity protocol is equal to the sum of their numbers of states, i.e., $O(\log^{3}n)$.
@@ 4 The Congruence problem
In this section, we show how our solution to the parity problem naturally extends to deciding congruences modulo any integer $m$. The main challenge is constructing a polynomial-time silent protocol for the general congruence problem. We chose not to adapt the succinct congruence protocol from Czerner et al. [19], as it remains unclear whether this protocol is silent. Font shape ‘T1/lmr/m/scit’ undefinedOur fast congruence protocol follows the same structure as in the parity case.
@@ 4.1 Slow, Stable and Silent Congruence Protocol
A formal description of the slow stable population congruence protocol SlowCongruence follows. This protocol uses $O(\log m)$ states and a crude analysis guarantees that it stabilises in expected time $O(n^{2})$. This protocol is used as a backup for the fast MC+ congruence protocol.
@@
4.2 Fast Congruence Protocol
The fast, silently stabilising congruence protocol can be constructed analogously to the fast parity protocol. The only remaining component of $m$-Congruence protocol (Algorithm 5) yet to be defined is the mechanism used to generate weights with masses $m,2m,4m,8m,\ldots$, referred to as $m$-WeightCreation (Algorithm 4). It differs from WeightCreation protocol only in a minor detail. To enable this new subprotocol, we consider the sequence of numbers $1=m_{0},m_{1},m_{2},\ldots,m_{k}=m$, which is the same as in the SlowCongruence protocol, and further extend this sequence with values $m_{k+i}=m2^{i}$, for $i>0$. And, for any integer $i>0$, we still get either $m_{i}=2m_{i-1}$ or $m_{i}=m_{i-1}+1$.
@@ 4.2 Fast Congruence Protocol
6:   Upon entering round $r$
7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $m_{k+r}=m2^{r}$).
8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$
9:   The leader transfers mass $m_{k+r}$ to the first free agent it encounters:
10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(k+r)$
11:   Distribute the mass by repeating for $i>0$:
12:   if $m_{i}=2m_{i-1}$ then $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$
13:   else $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(0)$
14:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
15:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
16:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
17:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
18:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
19:      it adopts state Max(r2)Max(r-2).
20:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
21:      it marks state $w_{r-3}$ as top.
22:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
23:   if a non-leader in state max(r2)max(r-2) enters round $r$
24:       interacting with state $\neq$ Max(r2)Max(r-2) then
25:      signal Anomaly
26:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
27:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
28:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
@@
5 Conclusion
In this work, we propose a new computing paradigm for population protocols, enabling the design of the first efficient, stable, and silent protocol for the parity and congruence problem. This new computing paradigm integrates several tools, including population weights, clocking mechanism, and efficient anomaly detection switching control to slow but always correct solutions. The new paradigm facilitates universal design of efficient multistage stable population protocols, enabling, e.g., efficient computation of more general congruencies. The power of the system of weights lies in the implicit conversion between unary and binary representations. In fact, due to this conversion, our parity protocol could also be amended to compute and report in a binary format (where, e.g., the consecutive bits of any integer value are stored in dedicated $O(\log n)$ agents) any subpopulation or even the whole population size, a problem posed in [21].
This raises a natural question: is binary representation of integers, both for input and output, better suited for designing efficient and stable population protocols? Although an efficient, stable multiplication protocol using binary representation is conceivable, achieving the same with unary representation is far more challenging, yet feasible (details omitted in this version). The primary obstacle is constructing a slow but stable protocol. Although recent years have seen remarkable progress in fast population protocols, advances in slow (polynomial time) yet stable protocols have been mainly limited to succinct protocols [19]. Our work indicates that further research is needed in this direction.

### v3

@@ 3.6 Efficient Parity Protocol – Proof of Theorem 1
By Theorem 2, the efficient MC+ parity protocol can be combined with a slow parity protocol, resulting in an efficient, stable parity protocol. This protocol is silent because the slow parity protocol is silent. The total number of rounds of the MC+ parity protocol is $R=O(\log^{2}n)$. By Lemma 2, the slow parity protocol requires polynomial time. By Theorem 2, the expected running time of the efficient stable parity protocol is $O(\log^{3}n)$. The total number of states in the combined parity protocol is equal to the sum of their numbers of states, i.e., $O(\log^{3}n)$.
@@ 4 The Congruence Problem
In this section, we argue that our parity protocol naturally extends to deciding congruences modulo any integer $m$. Our efficient congruence protocol (Section 4.2) follows the same structure as in the parity case. We do not adapt the polynomial‑time succinct protocol of Czerner et al. [19], as its silence is unclear. Instead, in Section 4.1, we present a protocol that is both stable and silent.
@@ 4.1 Slow, Stable and Silent Congruence Protocol
A formal description of the slow stable population congruence protocol SlowCongruence follows. This protocol uses $O(\log m)$ states and a crude analysis guarantees that it stabilises in expected time $O(n^{2})$. This protocol is used as a backup for the efficient MC+ congruence protocol.
@@
4.2 Efficient Congruence Protocol
The efficient, silently stabilising congruence protocol can be constructed analogously to the efficient parity protocol. The only remaining component of $m$-Congruence protocol (Algorithm 5) yet to be defined is the mechanism used to generate weights with masses $m,2m,4m,8m,\ldots$, referred to as $m$-WeightCreation (Algorithm 4). It differs from WeightCreation protocol only in a minor detail. To enable this new subprotocol, we consider the sequence of numbers $1=m_{0},m_{1},m_{2},\ldots,m_{k}=m$, which is the same as in the SlowCongruence protocol, and further extend this sequence with values $m_{k+i}=m2^{i}$, for $i>0$. And, for any integer $i>0$, we still get either $m_{i}=2m_{i-1}$ or $m_{i}=m_{i-1}+1$.
@@ 4.2 Efficient Congruence Protocol
6:   upon entering round $r$ interacting with an agent in state different than $m(2)$:
7:      if $r=0$ then
8:        the leader adopts state $\lambda_{0}$ and non-leaders adopt state $\phi$.
9:      else
10:        the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $m_{k+r}=m2^{r}$),
11:        a follower changes local state $w_{r-2}(0)$ to global state $w_{r-2}$ and local $w$,
12:        the leader changes state $\lambda_{r-1}$ to state $m(1)$,
13:        a follower changes state $w_{r-1}(i)$ to state $m(1)$, when $i>0$,
14:        the leader changes state $m(1)$ to state $m(2)$,
15:        a follower in state $m(1)$ signals Anomaly.
16:   upon entering round $r$ interacting with an agent in state $m(2)$:
17:      a follower adopts state $m(2)$.
18:   if interacting agents are in round $r$ then
19:       The leader transfers mass $m_{k+r}$ to the first free agent it encounters:
20:        $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(k+r)$
21:        Distribute the mass by repeating for $i>0$:
22:        if $m_{i}=2m_{i-1}$ then $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$
23:        else $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(0)$
24:        Epidemic of state $m(1)$: $m(1)+\text{\emph{any state}}\rightarrow m(1)+m(1)$.
25:  Compare $\bigcup W_{k}$ and $X$ using the Majority protocol
26:  (elements of $\bigcup W_{k}$ have global states $w_{k}$)
27:  if $|\bigcup W_{k}|<|X|$ then signal Anomaly
@@
Theorem 9.
5 Conclusions
In this work, we introduce the MC+ computing paradigm for population protocols, enabling, for the first time, efficient, stable, and silent protocols for parity and congruence. The paradigm combines three main components: population weights, a robust clocking mechanism, and effective anomaly detection supported by a fallback to a slow but always‑correct stable solution. This framework supports universally applicable multistage stable protocols, allowing, for instance, efficient computation of general congruences. The strength of the weight system lies in its implicit conversion between unary and binary representations. As a result, our parity protocol easily extends to compute and output (in binary) the size of any designated subpopulation, or even the entire population, with consecutive bits stored across $O(\log n)$ dedicated agents, thereby resolving a problem posed in [21].
This naturally raises the question of whether binary input and output offer a fundamentally better basis for designing efficient, stable population protocols than unary representation. While an efficient, stable multiplication protocol in binary seems plausible, achieving the same in unary remains far more difficult. Moreover, despite substantial recent progress on efficient population protocols, advances on slow yet stable protocols have largely been limited to succinct designs [16, 24, 19]. Our work indicates that this direction deserves much more attention. Finally, further refinements of the MC+ paradigm, especially more efficient stable solutions for leader election and majority computation, would be highly valuable.